OpenStudy (anonymous):

Find the volume of a sphere with diameter 15 feet. A. 1767.15 ft3 B. 994.02 ft3 C. 235.62 ft3 D. 471.24 ft3

OpenStudy (anonymous):

v=4/3 pie r^3 r=15/2

OpenStudy (anonymous):

i tried to solve it that way but it is wrong . can you show me the equation to get my answer ?

OpenStudy (anonymous):

@Aperogalics help !

OpenStudy (nubeer):

can you put up your steps.

OpenStudy (anonymous):

@alainabbyboo22 @03225186213 is correct absolutely

OpenStudy (anonymous):